# thanks: Corresponding author

### v1

@@ III.2 Classification results with spectrograms
Our classification test aims at distinguishing the elements of the dataset where a CCSN GW signal has been injected from those where only detector background noise is present. For a training of 25 epochs, the network found the minimum error in ${}^{\textrm{th}}$ epoch. The model obtained had an accuracy  [^1] The accuracy is defined as $\frac{TP+TN}{TP+TN+FP+FN}$, where TP (TN) corresponds to the correctly classified signals (background) and FP (FN) corresponds to the backgrounds (signals) that are misclassified. of 82%, providing the confusion matrix displayed in Figure 2. This figure shows that 92% of the background images are correctly classified, with a FPR of 8%. While the model only classifies 69% of the signals correctly, a true positive rate (TPR) of 90% is obtained. These results show that our model detects signals with high confidence, even if it only detects a modest amount of GW signals.
Figure 2: Confusion matrix for the classification test with ${}^{\textrm{th}}$ epoch weights showing the fractions of strains predicted as background noise or signal versus their actual labels.
With the previous training we further tested our model by performing additional studies with different datasets sorted by SNR and by distance to the source. Figure 3 (a) shows the relationship between the percentage of success of the network in detecting signals and the SNR of the injected GW. It is interesting to highlight that an accuracy of 95% and a precision of 93% is reached for SNR > 30. Although the network was trained to detect a GW generated at a distance of 20 kpc, it can also be used to infer GW signals at different distances. In Figure 3 (b), we observe that at shorter distances, the percentage of success is higher, as expected, decreasing as the source gets farther away. We observe a discrepancy for a distance of 20 kpc with respect to the value obtained previously during validation. This is due to the fact that when generating images for training and validation, those with SNR < 5 were eliminated, as mentioned above, a condition that we did not apply here.
@@ IV.1 Dataset generation
In our second approach, we apply DL techniques in the time domain. In this case we only use the 999 selected waveforms from the Richers et al. catalog, as described in Section II. For each of the selected waveforms we scale the strain $D\cdot h_{+}^{\textrm{opt}}$ to a distance of the source between 5 and 20 kpc (i.e. a CCSN that occurs in our galaxy). For the background noise we use public available data from both LIGO (Hanford and Livingston) and Virgo, with an initial GPS time $t_{\textrm{GPS}}=1253326755~{}\textrm{s}$, which corresponds to the end of September 2019, i.e. around the middle of the O3 run. We define a random time $t_{0}\in[6,1750]~{}\textrm{s}$ and select a segment of the noise with a window $[t_{0}-5,t_{0}+10]~{}\textrm{s}$ for the background sample. The waveforms were down-sampled to match the noise sampling frequency. In our analysis we used a sampling frequency of 4096 Hz, allowing faster performance without losing information. Each waveform was projected into each detector with random values for the angles defining the sky position (declination, right ascension, and polarization angle). Then, the projected waveforms are injected into noise samples for each detector, having the time of bounce of the signal ($t_{\textrm{b}}$) randomly set in the interval $[t_{0},t_{0}+0.80]~{}\textrm{s}$, for the specific case of datasets with one-second time window. For completeness, we also consider datasets of 0.50 s and 0.25 s with $t_{\textrm{b}}$ in the range $[t_{0},t_{0}+0.40]~{}\textrm{s}$ and $[t_{0},t_{0}+0.20]~{}\textrm{s}$, respectively.
@@
We perform whitening on the resulting strains (background and signal) through inverse spectrum truncation using the Amplitude Spectral Density (ASD) of the three detectors’ noise. We then apply a band-pass filter from 20 to 1000 Hz and notch filters at the individual frequencies of 60, 120, and 240 Hz, for the two LIGO strains, and 50, 100, and 200 Hz for the Virgo strains. We finally apply a crop on the strains to get the expected time window (1.00, 0.50, and 0.25 s).
@@
For the classification and parameter estimation tests in the time domain we use datasets of $10^{4}$ elements, where each element is composed of three time series, one for each detector. Examples of these time series are given in Figures 4 and 5, corresponding to one element of the signal and background samples, respectively. Table 1 characterizes the datasets used for classification and parameter estimation with time series. For the generation of the datasets, functions from the python [78] libraries SciPy [79], PyCBC [80], and GWpy [75], were used. We also used the libraries fast.ai [81] and tsai [82] for the classification and parameter estimation tests, described in the next sections.
@@
Figure 4: Example of a dataset’s element labeled as signal, with a waveform injected into noise from O3a, with a time window of 1 second (a) and 0.04 seconds (b).
@@
Figure 5: Example of a dataset’s element labeled as background, with only noise from O3a, with a time window of 1 second (a) and 0.04 seconds (b).
@@ IV.2 Deep Neural Network: architecture and methodology
Time-domain GW signals of CCSN are tested using a Residual Convolutional Neural Network (ResCNN), an implementation of CNN with residual learning blocks. This network, proposed in [83], has 6 convolutional layers and uses batch normalization techniques and different activation functions. The most used activation function is the Rectified Linear Unit (ReLU). Other versions of this activation function are also used in the architecture, i.e. Leaky ReLU (LReLU), Parametric ReLU (PReLU), and Exponential Linear Unit (ELU). It is desirable to use different activation functions in different layers to achieve better performance [83].
As the analysis of the time series needs the precise location of the signal features, the pooling operation was removed. Another reason for this choice is the fact that pooling reduces the number of inputs given to the next layer, limiting the information available. For this architecture, the usual Fully Connected Layer at the end of the network was replaced by a global average pooling and a final Softmax layer, so that the output of the network is a probability distribution of its own predictions. The architecture of this network is represented in Figure 6.
@@
We use the function fit_normal-_\__one_normal-_\__cycle to perform the training of the network, defining the maximum Learning Rate (LR) and the weight decay. This function of fastai internally uses the usual fit function but applies the 1cycle policy for better results [84]. The maximum LR and weight decay used should be adapted to each dataset and architecture [85]. In our case, to find the best values for our datasets we used another fastai function called lr_normal-_\__find based on the LR range test [86]. We verified that for our datasets the best results obtained were for weight decays and maximum LR of the order of $10^{-3}$.
For our parameter estimation computations, before giving any dataset to the network we perform a standard normalization of the signals’ parameters. To find the predictions of the network in both classification and parameter estimation, we use MC Dropout [87], passing each signal through the network 100 times. The prediction of the network is the mean value of all the 100 evaluations. Finally, we use PCA (from scikit-learn [88]) to evaluate the prediction distributions for the parameter estimation tests. PCA can be used for dimensionality reduction, being able to identify the most relevant information (features) in datasets. In this work, we only use the PCA function to perform Singular Value Decomposition (SVD) on the calibration plots, with predicted values versus true values of $D\cdot\Delta h$ and $f_{\textrm{peak}}$. We extract the two orthogonal eigenvectors to calculate the slope and with the singular values we obtain the standard deviation of the distributions, $\sigma$.
@@ IV.3 Classification results with time series
A classification test aims at obtaining a model that best allows to distinguish elements of a dataset where a CCSN GW signal has been injected from those of only background noise. Our best results are achieved for a 12-epoch training and considering a 0.50 threshold, which led to a 0.09 minimum validation loss. This model has an accuracy of 98%, providing the confusion matrix shown in Figure 7. Only 45 actual signals were predicted as background, which corresponds to 4.5% of all GW signals present in the validation set. In addition, only one element labeled as background was misclassified as a signal, giving a FPR of 0.10%. Figure 8 displays the receiver operating characteristic (ROC) curve of this classification model, with a 0.99 area under the curve (AUC).
@@
To better understand the results of the misclassified 45 signals (false negatives), Figure 9 shows the distribution of the model’s scores for each of those signals. In this plot we show our two key CCSN parameters, $D\cdot\Delta h$ (top panel) and $f_{\textrm{peak}}$ (bottom panel), as a function of the SNR. The different colors of the points represent the prediction scores given by the model. All misclassified signals have SNR $\leq$ 10, which is expected since in those cases it is difficult to distinguish the signal from the random peaks of the noise.
@@ IV.4 Parameter inference results with time series
A parameter inference test was performed to predict the values of the peak amplitude $D\cdot\Delta h$ (the difference between the highest and lowest points in the bounce signal normalized to the distance) and the peak frequency $f_{\textrm{peak}}$ (the highest frequency measured in the first 6 ms after bounce). As discussed above (see also [29]) these two parameters are the most relevant ones to characterize the waveform at bounce when $T/|W|<0.06$, as $D\cdot\Delta h\propto T/|W|$ and $f_{\textrm{peak}}\propto\sqrt{\rho_{c}}$ for a large group of EoS [20].
Three datasets of $10^{4}$ elements each were analyzed, for different choices of time windows, i.e. 1.00, 0.50, and 0.25 s, as mentioned previously. For all the parameter estimation tests, we trained the network for 30 epochs setting the LR to $2\times 10^{-3}$ and the weight decay to $10^{-3}$. As expected, there is a relation between the size of the time window used and the training time needed for each epoch. For the 1.00 s dataset, each epoch took around 90 s to complete. By decreasing the window to 0.50 s, the time needed for each epoch also decreased to around 50 s. And lastly, for the 0.25 s dataset, each epoch took around 27 s to complete. As we decrease the time window of the dataset, the training time also decreases, almost in a linear way.
@@
Figure 11: Calibration plots of predicted vs true $D\cdot\Delta h$ (left) and predicted vs true $f_{\textrm{peak}}$ (right) with SNR $\geq 15$ for the different datasets. As in Fig. 10 the white dashed line represents the ideal case (slope 1) and the white dotted line represents the slope of the distribution calculated with the PCA function.
@@
It is worth noticing that horizontal scattered lines are visible in Figure 10 for all the distributions of the calibration plots. In order to understand this feature we generated new datasets with a minimum SNR of 15, consistent with other studies which place constraints on the minimum SNR estimated to detect GWs from CCSN [14]. The new results for the parameter estimation of $D\cdot\Delta h$ and $f_{\textrm{peak}}$ on the three different datasets are shown in Figure 11. With this choice of minimum SNR the horizontal scattered lines disappear. We also see an improvement in the values of the slopes, with the best results corresponding to the 0.25 seconds dataset. In this case, for the $D\cdot\Delta h$ distribution, the slope is 0.884 and the standard deviation is $\sigma_{D\cdot\Delta h}=52.6$ cm. For the $f_{\textrm{peak}}$ distribution, the results yield a slope of 0.917 and a standard deviation of $\sigma_{f_{\textrm{peak}}}=18.3$ Hz.
@@
The probability distributions of the difference between true and predicted values of $D\cdot\Delta h$ and $f_{\textrm{peak}}$ are shown in the two plots of Figure 12, for the 0.25 seconds dataset with minimum SNR of 15. For $D\cdot\Delta h$, a mean value of $\mu_{[D\cdot\Delta h_{\textrm{true}}-D\cdot\Delta h_{\textrm{pred.}}]}=0.5~{}\pm~{}76.6$ cm is obtained. As for the $f_{\textrm{peak}}$ case, the mean value obtained is $\mu_{[f_{\textrm{peak, true}}-f_{\textrm{peak, pred.}}]}=~{}-0.2~{}\pm~{}26.4$ (Hz). We can clearly see that the inference of $f_{\textrm{peak}}$ is much more accurate than $D\cdot\Delta h$. Taking into account that in our datasets $D\cdot\Delta h\in[100,850]$ cm and $f_{\textrm{peak}}\in[650,950]$ Hz, we can consider the mean values and standard deviations obtained quite reasonable.
@@ IV.5 Testing the network with a different catalog
Table 2: Parameter estimation results for the 30 selected waveforms from the Mitra et al. catalog [32, 89]. The models are grouped by the degree of differential rotation $A$, corresponding, from top to bottom to 300 km (A1), 634 km (A3), 1268 km (A4), and 10000 km (A5). For each waveform we report the prediction, for both $D\cdot\Delta h$ and $f_{\textrm{peak}}$, in the form $\mu\pm\sigma$, and the true value in brackets. When the true values are within $1\sigma$ of the predictions they are written in boldface.
We selected 30 waveforms of this new catalog and, for each one of them, generated a dataset with 100 elements as we did before. We fixed the time window to 0.25 s and the minimum SNR to 15, which corresponds to the conditions that led to our best model. Moreover, we allowed the model to vary its response by using MCDropout. By performing signal injections in different noise conditions as well as changes in the model response, we obtain a more realistic implementation of a true GW detection. For a given dataset, prediction values were determined for each element and, with the collection of the 100 predicted values, the mean value, $\mu$, and the standard deviation, $\sigma$, were calculated for both parameters, $D\cdot\Delta h$ and $f_{\textrm{peak}}$.

### v2

@@ III.2 Classification results with spectrograms
Our classification test aims at distinguishing the dataset elements where a CCSN GW signal has been injected from those where only detector background noise is present. For a training of 25 epochs, the network found the minimum error in the 13th epoch. The model obtained had an accuracy [^1] The accuracy is defined as $\frac{TP+TN}{TP+TN+FP+FN}$, where $TP$ ($TN$) corresponds to the correctly classified signals (background) and $FP$ ($FN$) corresponds to the backgrounds (signals) that are misclassified. of 82%, providing the confusion matrix displayed in Fig. 2. This figure shows that 92% of the background images are correctly classified, with a false positive rate (FPR) of 8%. While the model only classifies 69% of the signals correctly, a precision of 90% is obtained. These results show that our model detects signals with high confidence, even if it only detects a modest amount of GW signals.
Figure 2: Confusion matrix for the classification test with the 13th epoch weights showing the fractions of strains predicted as background noise or signal versus their actual labels.
III.3 Testing the classification model
We further tested our model by performing additional studies with two different datasets sorted by SNR and by distance to the source. One dataset consists of 1500 images arranged in groups differentiated by five SNR units, with fixed distance and sky position, similar to the datasets used for training. The other dataset consists of 4000 images with distances ranging between 5 and 40 kpc, keeping a fixed sky position and without any limits on the SNR. Figure 3(a) shows the relationship between the percentage of success of the network in detecting signals and the SNR of the injected GW. It is interesting to highlight that an accuracy of 95% and a precision of 93% is reached for SNR > 30. Although the network was trained to detect a GW generated at a distance of 20 kpc, it can also be used to infer GW signals at different distances. In Fig. 3 (b), we observe that at shorter distances, the percentage of success is higher, as expected, decreasing as the source gets farther away. We observe a discrepancy for a distance of 20 kpc with respect to the value obtained previously during validation. This is due to the fact that when generating images for training and validation, those with SNR < 5 were eliminated, as mentioned above, a condition that we did not apply here.
@@ IV.1 Dataset generation
In our second method, we apply DL techniques in the time domain. While significant patterns may be more easily recognized in the time-frequency representation, some information might be lost during the Q-transform process, which would not occur when using time series directly. In this case, we only use the 999 selected waveforms from the Richers et al. catalog, described in Sec. II. For each of the selected waveforms, we scale the strain $D\cdot h_{+}^{\textrm{opt}}$ to a randomly chosen distance of the source between 5 and 20 kpc (i.e., a CCSN that occurs in our Galaxy). For the background noise we use publicly available data from both LIGO (Hanford and Livingston) and Virgo, with an initial Global Positioning System (GPS) time $t_{\textrm{GPS}}=1253326755~{}\textrm{s}$, which corresponds to the end of September 2019, i.e., around the middle of the O3 run. We define a random time $t_{0}\in[6,1750]~{}\textrm{s}$ and select a segment of the noise time series with a window $[t_{0}-5,t_{0}+10]~{}\textrm{s}$ for the background sample. The waveforms were down-sampled to match the noise sampling frequency. In this analysis we used a sampling frequency of 4096 Hz, allowing faster performance without losing information. Each waveform was projected into each detector with random values for the angles defining the sky position (declination, right ascension, and polarization angle). Then, the projected waveforms are injected into noise samples for each detector, having the time of bounce of the signal ($t_{\textrm{b}}$) randomly set in the interval $[t_{0},t_{0}+0.80]~{}\textrm{s}$, for the specific case of datasets with one-second time window. For completeness, we also consider datasets of 0.50 s and 0.25 s with $t_{\textrm{b}}$ in the range $[t_{0},t_{0}+0.40]~{}\textrm{s}$ and $[t_{0},t_{0}+0.20]~{}\textrm{s}$, respectively.
@@
We perform whitening on the resulting strains (both for the background and signal classes) through inverse spectrum truncation using the amplitude spectral density of the three detectors’ noise. We then apply a bandpass filter from 20 to 1000 Hz and notch filters at the individual frequencies of 60, 120, and 240 Hz for the two LIGO strains and 50, 100, and 200 Hz for the Virgo strains. We finally apply a crop on the strains to get the expected time window (1.00, 0.50, and 0.25 s). At this point, a selection is made on the resulting whitened background section to exclude high-intensity glitches (such as blips). After whitening, the maximum absolute amplitude is around $1.5\times 10^{-21}$, so we only allow segments with a maximum absolute amplitude below $2.7\times 10^{-21}$.
@@
For the classification and parameter estimation tests in the time domain, we use datasets of $10^{4}$ elements, where each element is composed of three time series, one for each detector. Examples of these time series are given in Figs. 4 and 5, corresponding to one element of the signal and background samples, respectively. Table 1 characterizes the datasets used for classification and parameter estimation with time series. For the generation of the datasets, functions from the PYTHON [78] libraries SciPy [79], PyCBC [80], and GWpy [75] were used. We also used the libraries FAST.AI [81] and TSAI [82] for the classification and parameter estimation tests, described in the next sections.
@@
Figure 4: Example of a dataset’s element labeled as signal, with a waveform injected into noise from O3a, with a time window of 1 s (a) and 0.04 s (b).
@@
Figure 5: Example of a dataset’s element labeled as background, with only noise from O3a, with a time window of 1 s (a) and 0.04 s (b).
@@ IV.2 Deep neural network: Architecture and methodology
Time-domain GW signals of CCSN events are tested, both for classification and parameter estimation, using a residual convolutional neural network (ResCNN). This network is an implementation of the CNN with residual learning blocks, proposed in [65]. It has six convolutional layers and uses batch normalization techniques and different activation functions. The most used activation function is the rectified linear unit (ReLU). Other versions of this activation function are also used in the architecture, i.e., leaky ReLU (LReLU), parametric ReLU (PReLU), and exponential linear unit (ELU). It is desirable to use different activation functions in different layers to achieve better performance [65]. As the analysis of the time series needs the precise location of the signal features, the pooling operation was removed. Another reason for this choice is the fact that pooling reduces the number of inputs given to the next layer, limiting the information available. For this architecture, the usual fully connected layer at the end of the network was replaced by a global average pooling and a final softmax layer, so that the output of the network is a probability distribution of its own predictions. The architecture of this network is represented in Fig. 6.
@@
We use the function fit__\__one__\__cycle to perform the training of the network, defining the maximum learning rate (LR) and the weight decay. This function of FAST.AI internally uses the usual fit function but applies the 1cycle policy for better results [83]. The maximum LR and weight decay used should be adapted to each dataset and architecture [84]. In our case, to find the best values for our datasets we used another FAST.AI function called lr__\__find based on the LR range test [85]. We verified that for our datasets the best results obtained were for weight decays and maximum LR of the order of $10^{-3}$.
For our parameter estimation computations, before giving any dataset to the network, we perform a standard normalization of the signals’ parameters. To find the predictions of the network in both classification and parameter estimation, we use Monte Carlo (MC) dropout [86], passing each signal through the network 100 times. The prediction of the network is the mean value of all the 100 evaluations. Finally, we apply PYTHON’S PCA function from SCIKIT-LEARN [87] on the two-dimensional distributions of predicted versus true values, obtained on the parameter estimation tests, according to what follows. Although PCA is typically used for dimensionality reduction by identifying the most relevant features in datasets, we do not use this capability of PCA here. Instead, we use the PYTHON function solely to perform singular value decomposition (SVD) on the calibration plots, comparing predicted values to true values of $D\cdot\Delta h$ and $f_{\textrm{peak}}$. We extract the two orthogonal eigenvectors to calculate the slope, and with the singular values we obtain the standard deviation of the distributions, $\sigma$.
@@ IV.3 Classification results with time series
A classification test aims at obtaining a model that best allows to distinguish elements of a dataset where a CCSN GW signal has been injected from those of only background noise. Our best results are achieved for a 12-epoch training and considering a 0.50 threshold, which led to a 0.09 minimum validation loss. This model has an accuracy of 98%, providing the confusion matrix shown in Fig. 7. Only 45 actual signals were predicted as background, which corresponds to 4.5% of all GW signals present in the validation set. In addition, only one element labeled as background was misclassified as a signal, giving a FPR of 0.10%. Figure 8 displays the receiver operating characteristic (ROC) curve of this classification model, with a 0.99 area under the curve.
@@
To better understand the results of the misclassified 45 signals (false negatives), Fig. 9 shows the distribution of the model’s scores for each of those signals. In this plot we show our two key CCSN parameters, $D\cdot\Delta h$ (top panel) and $f_{\textrm{peak}}$ (bottom panel), as a function of the SNR. The different colors of the points represent the prediction scores given by the model. All misclassified signals have SNR $\leq$ 10, which is expected since in those cases it is difficult to distinguish the signal from the random peaks of the noise.
@@
For completeness we have built a spectrograms dataset corresponding to the O3a data used here, to test the network applied in Sec. III.2 to spectrograms. As expected, when changing the background noise of the spectrograms to O3a, the accuracy of the network goes up to 95%, reflecting the better noise conditions of O3a when compared to O2.
@@ IV.4 Parameter inference results with time series
A parameter inference test was performed to predict the values of the GW strain amplitude ($\Delta h$) multiplied by the luminosity distance ($D$), $D\cdot\Delta h$ (the difference between the highest and lowest points in the bounce signal normalized to the distance) and the peak frequency, $f_{\textrm{peak}}$ (the highest frequency measured in the first 6 ms after bounce). As discussed above (see also [29]) these two parameters are the most relevant ones to characterize the waveform at bounce when $T/|W|<0.06$, as $D\cdot\Delta h\propto T/|W|$ and $f_{\textrm{peak}}\propto\sqrt{\rho_{c}}$ for a large group of equations of state [20].
Three datasets of $10^{4}$ elements each were analyzed for different choices of time windows, i.e. 1.00, 0.50, and 0.25 s, as mentioned previously. For all the parameter estimation tests, we trained the network for 30 epochs setting the LR to $2\times 10^{-3}$ and the weight decay to $10^{-3}$. As expected, there is a relation between the size of the time window used and the training time needed for each epoch. For the 1.00 s dataset, each epoch took around 90 s to complete. By decreasing the window to 0.50 s, the time needed for each epoch also decreased to around 50 s. And lastly, for the 0.25 s dataset, each epoch took around 27 s to complete. As we decrease the time window of the dataset, the training time also decreases, almost in a linear way.
@@
Figure 11: Calibration plots of predicted vs true $D\cdot\Delta h$ (left) and predicted vs true $f_{\textrm{peak}}$ (right) with SNR $\geq 15$ for the different datasets. As in Fig. 10 the white dashed line represents the ideal case (slope 1), and the white dotted line represents the slope of the distribution calculated with the PCA function.
@@
It is worth noticing that horizontal scattered lines are visible in Fig. 10 for all the distributions of the calibration plots. In order to understand this feature we generated new datasets with a minimum SNR of 15, consistent with other studies which place constraints on the minimum SNR estimated to detect GW signals from CCSN events [14]. The new results for the parameter estimation of $D\cdot\Delta h$ and $f_{\textrm{peak}}$ on the three different datasets are shown in Fig. 11. With this choice of minimum SNR the horizontal scattered lines disappear. We also see an improvement in the values of the slopes, with the best results corresponding to the 0.25 seconds dataset. In this case, for the $D\cdot\Delta h$ distribution, the slope is $0.884\pm 0.008$, and the standard deviation is $\sigma_{D\cdot\Delta h}=(52.6\pm 10.7)$ cm. For the $f_{\textrm{peak}}$ distribution, the results yield a slope of $0.917\pm 0.006$ and a standard deviation of $\sigma_{f_{\textrm{peak}}}=(18.3\pm 3.9)$ Hz.
@@
The probability distributions of the difference between true and predicted values of $D\cdot\Delta h$ and $f_{\textrm{peak}}$ are shown in the two plots of Fig. 12 for the 0.25 seconds dataset with minimum SNR of 15. For $D\cdot\Delta h$, a mean value of $\mu_{[D\cdot\Delta h_{\textrm{true}}-D\cdot\Delta h_{\textrm{pred.}}]}=0.5~{}\pm~{}76.6$ cm is obtained. As for the $f_{\textrm{peak}}$ case, the mean value obtained is $\mu_{[f_{\textrm{peak, true}}-f_{\textrm{peak, pred.}}]}=~{}-0.2~{}\pm~{}26.4$ (Hz). We can clearly see that the inference of $f_{\textrm{peak}}$ is much more accurate than $D\cdot\Delta h$. Taking into account that in our datasets $D\cdot\Delta h\in[100,850]$ cm and $f_{\textrm{peak}}\in[650,950]$ Hz, we can consider the mean values and standard deviations obtained quite reasonable.
@@ IV.5 Testing the network with a different catalog
Table 2: Parameter estimation results for the 30 selected waveforms from the Mitra et al. catalog [32, 88]. The models are grouped by the degree of differential rotation $A$, corresponding from top to bottom to 300 km (A1), 634 km (A3), 1268 km (A4), and 10000 km (A5). For each waveform we report the prediction, for both $D\cdot\Delta h$ and $f_{\textrm{peak}}$, in the form $\mu\pm\sigma$ and the true value in brackets. When the true values are within $1\sigma$ of the predictions they are written in boldface.
We selected 30 waveforms of this new catalog and, for each one of them, generated a dataset with 100 elements as we did before. We fixed the time window to 0.25 s and the minimum SNR to 15, which corresponds to the conditions that led to our best model. Moreover, we allowed the model to vary its response by using MC dropout. By performing signal injections in different noise conditions as well as changes in the model response, we obtain a more realistic implementation of a true GW detection. For a given dataset, prediction values were determined for each element and, with the collection of the 100 predicted values, the mean value $\mu$ and the standard deviation $\sigma$ were calculated for both parameters, $D\cdot\Delta h$ and $f_{\textrm{peak}}$.